OpenStudy (anonymous):

when the cost of a car is multiplied by 0.09 the result is $1780. find the cost of the car.

OpenStudy (anonymous):

let its price be x 0.09x=1780 find x

OpenStudy (anonymous):

can you find x?

OpenStudy (shamim):

19777.7777

OpenStudy (anonymous):

try and write 0.09 as 9/100

OpenStudy (anonymous):

i dont understand howto find x. it doesnt make sense

OpenStudy (shamim):

ok u can divide 1780 by 0.09. u will get the value of ur car or x

OpenStudy (anonymous):

0.09x = 9x/100 = 1780. Multiply both sides by 100/9

OpenStudy (anonymous):

\[ \frac{9x}{100} \times \frac{100}{9} = 1780 \] \[ \Rightarrow x = \frac{1780 \times 100}{9} = \frac{178000}{9} = 19777.\bar{7} \]